OpenStudy (anonymous):

if t=(-1,3) and r= (2,-4), what is the vector which represents t-(1/2)r?

OpenStudy (anonymous):

its (-2,5) right ?

OpenStudy (anonymous):

(1/2)r = vector (1, -2) so t-(1/2)r = ([-1-1],[3 --2]) = ... Vector (-2,5)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You are welcome.